Smart Robot Company is a company that makes very small robots. The company's October 2020 budget is shown below: Price per unit: $300 Variable expenses per unit: $200 Fixed expenses for October: $75,000 Unit sales in October: 10,000 (1) Calculate the Margin of Safety (2) What is the margin of safety as a percentage of sales? Please show all work for the problem.
Answer #1

Contribution margin=Sales-Variable cost  

=(300-200)=$100 per unit

Breakeven=Fixed expenses/Contribution margin

=(75000/100)=750 units

Margin of Safety=Total sales-Breakeven sales

=10,000-750

=9250 units

=(9250*300)=$2775000

Margin of Safety %=Margin of Safety/Total sales

=9250/10,000

=92.5%